Write the first 4 terms of the given series, furthermore, determine if the series converges or diverges using the Theorem: Condition necessary for convergence.

In () k 3k + 1 k=1
Answer #1

Solh: Given that 00 In 15 E kal E ak 3kti ke1 First your teams are a = ln (100+1 ti) Inlt) as= ln (12/2/22 +1 :) = In (&). 23